OpenStudy (anonymous):

Change 4x – 2y = 9 into slope-intercept form

OpenStudy (anonymous):

2y = 4x - 9 y = 4/2x - 9/2 y = 2x -9/2 Slope = 2 Y-intercept = -9/2

Directrix (directrix):

m = 2 y-intercept is -9/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

4x – 2y = 9 2y = 4x - 9 ->y = 2x - 9/2

OpenStudy (anonymous):

why would you waste time answering a question that's already answered?

OpenStudy (anonymous):

I never look at what the others doing as you are :P

OpenStudy (anonymous):

but then you're just doing all the work for nothing... you should check other people's work before submitting your own